OpenStudy (anonymous):

Simplify \({\sqrt(2)-\sqrt(6)}/{\sqrt(2)+\sqrt(6)}\)

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{ \sqrt{2} - \sqrt{6} }{ \sqrt{2} + \sqrt{6} } }\]

OpenStudy (solomonzelman):

Multiply top and bottom by conjugate, or by \[\sqrt{2} - \sqrt{6}\]

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{ \sqrt{2} - \sqrt{6} }{ \sqrt{2} + \sqrt{6} } } \times \huge\color{red}{ \frac{ \sqrt{2} - \sqrt{6} }{ \sqrt{2} - \sqrt{6} } }\]

OpenStudy (anonymous):

Ok i see that but how to I multiple them together?

OpenStudy (anonymous):

I never multiplied squareroots

OpenStudy (solomonzelman):

\[\huge\color{green}{ \frac{ (\sqrt{2} - \sqrt{6})( \sqrt{2} - \sqrt{6}) }{ (\sqrt{2} + \sqrt{6} ) (\sqrt{2}- \sqrt{6})} }\]

OpenStudy (solomonzelman):

\[\huge\color{red}{ \sqrt{x} \times \sqrt{x} =x }\]

OpenStudy (solomonzelman):

On the bottom, the square root are going to cancel.

OpenStudy (anonymous):

oh so would the top be 2 -6?

OpenStudy (solomonzelman):

Close, on the top there is also a middle term.

OpenStudy (solomonzelman):

\[\huge\color{red}{ (a-b)^2=a^2-2ab+b^2 }\]use this formula to do the top.

OpenStudy (anonymous):

\[(2 - 6)^2 = 2^2 - 2(2)(-6)+(-6)^2\]

OpenStudy (anonymous):

64?

OpenStudy (solomonzelman):

Well, it's \[( \sqrt {2}-\sqrt {6} )^2\]

OpenStudy (solomonzelman):

And not \[(2-6)^2\]

OpenStudy (anonymous):

oh how would we do it with the squareroots?

OpenStudy (solomonzelman):

\[(\sqrt{2}-\sqrt{6})^2=\sqrt{2}~^2~~-~~2\times \sqrt{2} \times \sqrt{6}~~-~~(-\sqrt{6})~^2\]

OpenStudy (solomonzelman):

\[2-2\sqrt{12}+6\]\[8-2\sqrt{12}\]\[8-4\sqrt{3}\]

OpenStudy (anonymous):

ok i get that now but how would we do the bottom you said it cancels out

OpenStudy (solomonzelman):

Sorry, I got disconnected.

OpenStudy (solomonzelman):

\[(a-b)(a+b)=a^2-b^2\]\[(\sqrt{2}-\sqrt{6})~(\sqrt{2}+\sqrt{6})~=~\sqrt{2}~^2~~~-~~~\sqrt{6}~^2=~2~-~6~~=~~-4\]

OpenStudy (anonymous):

\(-2\sqrt{3}\)?

OpenStudy (solomonzelman):

No, it's not, \[-2\sqrt{3}\]

OpenStudy (solomonzelman):

Do you see how I did the bottom?

OpenStudy (anonymous):

oh wait so you got -4 for the bottom? is see that somewhat

OpenStudy (solomonzelman):

Yes, -4 on the bottom.

OpenStudy (solomonzelman):

So so far we have \[\huge\color{blue}{ \frac{8 - 4\sqrt{3}}{-4} }\]

OpenStudy (solomonzelman):

Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8-4=2 so \(2\sqrt3\) the top

OpenStudy (solomonzelman):

You can't subtract those, they are not like terms. One is \[8~~~~~(whole~~number)\] and the other is \[4\sqrt{3}~~~~~~~~(radical)\]

OpenStudy (anonymous):

oh so how would I simplify those?

OpenStudy (solomonzelman):

factor the top out of -4. \[\huge\color{blue}{ \frac{8 - 4\sqrt{3}}{-4} }\] \[\huge\color{blue}{ \frac{-4(-2 +1\sqrt{3})}{-4} }\] \[\huge\color{blue}{ \frac{-4(-2 +\sqrt{3})}{-4} }\] \[\huge\color{blue}{ (-2 +\sqrt{3}) }\]

OpenStudy (anonymous):

how did you factor out 8-4sqrt(3) -4*-2 = 8 -4*1 = -4 oh never mind got it thank you so much :D

OpenStudy (solomonzelman):

Are you sure you got it?

OpenStudy (anonymous):

yup the -4 cancels the -4 on the bottom leaving us with -2srt(3)

OpenStudy (solomonzelman):

well, not exactly, it's \[-2 + \sqrt{3}\]we didn't add, just factored out of -4.

OpenStudy (anonymous):

oh okay I understand now :) thanks

OpenStudy (solomonzelman):

YW!